Is the function y = x / (x 2 - 3) EVEN, ODD or neither
Answer by stanbon(75887)

Is the function y = x / (x^2 - 3) EVEN, ODD or neither
--------------
Compare f(x), f(-x), and -f(-x)
f(x) = x/(x^2-3)
f(-x) = -x/[(-x)^2-3] = -x/[x^2-3]
-f(-x) = x/[x^2-3]
-------------------------
Notice that f(x) is not equal to f(-x) so f is not EVEN
Notice that f(x) is equal to -f(-x) so f is ODD
==============
Cheers,
Stan H.